An accelerator at a Physics Laboratory can accelerate protons to a Kinetic energy of 10 ^ 12 * eV The rest mass of a proton is m * c ^ 2 = 9.38 * 10 ^ 6 * cV . What is the speed of that proton?


1
Expert's answer
2022-05-30T15:16:50-0400
"E_k=\frac{mc^2}{\sqrt{1-v^2/c^2}}-mc^2""10^{12} =\frac{9.38 * 10 ^ 6}{ \sqrt{1-v^2/c^2}}-9.38 * 10 ^ 6"

"1-v^2\/c^2=1\/((E_k\/mc^2+1)^2"

"v=c\sqrt{1-1/(E_k/mc^2+1)^2}"

"v=c\sqrt{1-1/(10 ^{ 12}/9.38*10^{6}+1)^2}=c"

impossible